What is the perimeter of the trapezoid?



1
Expert's answer
2022-03-24T05:26:59-0400

The formula is P=T+B+L+R, where P equals the perimeter of the trapezoid, and the variables T equals the length of the top base of the trapezoid, B equals the length of the bottom base, L equals the length of the left side, and R equals the length of the right side.

  • For example, if you have a trapezoid with a top base of 4 cm, a bottom base of 5 cm, and two side lengths of 2 cm, your formula will look like this:
  • P=4+5+2+2

Add the side lengths together. This will give you the perimeter of your trapezoid.

  • P=4+5+2+2
  • P=13
  • So, the perimeter of the trapezoid is 13 cm.